Fermi National Laboratory


Accelerator Update

Friday January 30
The day shift began with the with D-zero (D0) in access to repair a detector water leak, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 10:40 AM, a Linac RF station (KRF1) tripped off. It reset okay.

A problem at 11:48 AM that showed beam being sent to Meson when there was no beam shut off all beam for about a half hour. This ended up being a noise problem and was fixed.

A Booster power supply (EAPS) tripped off at 1:10 PM. An expert investigated and said that a Booster RF station (BRF7) went down hard and caused the power supply to trip. RF technicians found and fixed a broken relay. They said the original trip was due to a TLG problem.

At 1:10 PM, operators made the Tevatron (TeV) ready for an access. Maintenance techs were out of the TeV by 3:41 PM.

The harmonic filter at the Master Substation would not close at 4:10 PM. At 4:30 PM, the High Voltage expert said that the filter would need repairs; feeder 23 could not be racked in at the end of the TeV access. Techs had these repairs completed by 7 PM.

At 7:49 PM, D0 experimenters reported that they found a silicon detector power supply problem that would take four to five hours to repair.

D0 was out of access at 10 PM, but they continued with power supply work.

Operations put the TeV into shot setup at 11:58 PM.

Saturday January 31
The midnight shift began with the TeV in shot setup and with MiniBooNE taking beam.

At 2:25 AM, Operations established store 3210 with an initial luminosity of 48.76E30. Pbar resumed stacking.

MiniBooNE tripped off at 7:57 AM. It reset okay.

Switchyard experts began a study period at 12:46 PM.

Operations terminated store 3210 at 4:56 PM.

Techs completed their D0 power supply work at 6:37 PM. Operations prepared D0 for an access to close up their detector. They were out of access by 8:33 PM.

At 8:56 PM, Operations put the TeV through a dry squeeze.

Operations put the TeV into shot setup at 9:23 PM.

The Main Injector (MI) suffered a vacuum burst in its RF section at 9:32 PM. The vacuum pumps all reset okay.

At 10:14 PM, a Linac RF station (KRF3) tripped off for the first of four times in a forty-five minute time period. A Linac expert advised operators on how to adjust a reference voltage. The operators had the station operating normally by 11:01 PM.

Sunday February 1
The midnight shift began with the TeV in shot setup and with MiniBooNE taking beam.

At 12:53 AM, Operations established store 3212 with an initial luminosity of 51.2E30. Pbar resumed stacking.

At 8:01 AM, a MI RF station (MIRF14) tripped off due to a power supply grid problem.

A Switchyard expert began a nine-hour study period at 10:11 AM.

MiniBooNE tripped off at 11:36 AM due to a power supply problem (E:HV865). Experts had the supply repaired and MiniBooNE back on by 1:13 PM.

At 3:35 PM, a TeV RF station (TRF8) tripped off due to a heater ground fault. An expert said that temperature was really the problem and should normalize soon and then the station could be turned back on.

The Pbar dump cooling system tripped off at 5:36 PM. Operations stopped stacking until 6:23 PM when an expert got the system working again.

Monday February 2
The midnight shift began with Operations monitoring store 3212, with Pbar stacking, and with MiniBooNE taking beam.

At 12:50 AM, a MI RF station (MIRF13) tripped off for the first of seven times over four and a half hours.

A TeV expert came in at 6 AM to conduct an end-of-store study.

A Recycler expert began a study period at 6:47 AM.

Today's Plans
The plans for today are to end the store shortly after 8 AM and then go into a new shot setup.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/2/2004   email Fermilab